OpenStudy (anonymous):

A particle moves on a line away from its initial position so that after t hours it is s = 4t^2 + t miles from its initial position. Find the average velocity of the particle over the interval [1, 4]. Include units in your answer.

OpenStudy (anonymous):

@Luigi0210 :)

OpenStudy (luigi0210):

Yea, I can't do word problems, sorry bud P: @welshfella looks like they've got something tho :D

OpenStudy (welshfella):

calculate s for t = 1 and t = 4 then find the difference to get the distance travelled in that period of 3 seconds average velocity = distance / time

OpenStudy (welshfella):

time will equal 3 secnds

OpenStudy (welshfella):

sorry 3 hours

OpenStudy (anonymous):

can u please show me the steps? :)

OpenStudy (anonymous):

ok so i would do (t=1) s=5 (t=4) s=68 ? :)

OpenStudy (welshfella):

find s for t = 4: s = 4*(4)^2 + 4 = ? then do the same for t = 1 then subtract the first calculation from the second finally divide the result by 3

OpenStudy (anonymous):

Do you know how to calcualte the average rate of change?

OpenStudy (anonymous):

Use this formula: AROC is: F(4) - F(1) / (4 - 1)

OpenStudy (anonymous):

alright so its @welshfella 68-5=63/3 = 21? :D

OpenStudy (welshfella):

thats it

OpenStudy (anonymous):

@Fzxt , okay so its 68-5/3 :) yes

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

So the average velocity is...how do i do the units i think its: 21 mph?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Yep, the units should be mentioned in whatever question you are doing.

OpenStudy (anonymous):

THANK U BOTH!!!!!!!

OpenStudy (anonymous):

do you guys have time for 5 more? :)

OpenStudy (anonymous):

You should make a seperate question.

OpenStudy (anonymous):

yeah I will :) thank u for the help on this 1!!!

OpenStudy (anonymous):

No problemo.

OpenStudy (anonymous):

idk who to choose as best response :P lol